A fan rotating at 3 rev/s is switched off what is it's angular acceleration in rev/² when it stops?
1
Expert's answer
2020-06-29T14:05:56-0400
"\alpha=\frac{\omega}{t}"

"\alpha=\frac{3}{18}=0.17\frac{rev}{s^2}"